OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 33 feet and the child ride has a radius of 22 feet, what is the ratio between the volumes of each ride?

OpenStudy (ganpat):

Volume of Cone = 1/3 * pi * r square * h Volume adult/ Volume childrens = (1/3 * pi * r square * h) adult / (1/3 * pi * r square * h) children = ( 33 * 33)/ (22 * 22) = 9/4.. assuming for same height.. and rest all terms get cancelled.. understood ??

OpenStudy (anonymous):

yes thanks